KTeV Shift Summary - EVE 11/27/99 (Sat.)

Earl C. Swallow (earls@elmhurst.edu)
Sat, 27 Nov 1999 23:31:39 -0600 (CST)


Shift Summary for EVE shift, 11/27/99 (Sat.)

On Shift: Ashkan Alavi-Harati and Earl Swallow

Hours of beam available at all intensities: 8.0 (out of 8.5)
Hours of beam available at nominal intensity: 8.0
Hours of E799 data (special runs not included): 6.9
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time & intensity:

DAQ problems & NM4 interlocks drop plus energy conservation, a few
aborts, and KFRENDs.

========================================================================
Runs ended this shift:

15069 -- BEAM_799 (15 M triggers)

Tapes brought to FCC this shift:

UPA,B,C199-201, UPS067 -- Runs 15065, 15066, 15068
UPA,B,C202-204, UPS068 -- Run 15069

========================================================================
Accesses during shift:

None.

========================================================================
Overview of shift:

Smooth data taking until DAQ problems arise.

========================================================================
Description of Shift:

15069 15:30 EVEning Shift Begins -- smooth running in progress.

15:47 Online Split requests UPF051 ---> KTEV4/tps3d2.

16:00 Online Split requests UPE014 ---> KTEV2/tps3d2.
[UPK026 removed from drive before mount.]

19:13 Online Split requests UPX015 ---> KTEV3/tps2d3.

20:15 DC4X-H-R problem is back again after a week of proper
operation. AA investigates and identifies downstairs
patch panels or transformers as candidates for
problem source.

20:45 Online Split requests UPT016 ---> KTEV1/tps3d2.

15069 21:00 Run 15069 aborted because Taku says he is full --
however, tapes are really only a bit under 1/2 full!
We end run anyway (and mount new tapes) to guard
against any data loss.

21:20 Trigger refuses to start for new run, and
drt_fresh_start doesn't help. Leo and AA mount
attack on problem.

22:06 NM4 interlocks drop -- ECS joins operators to reset.
There may have been a dead fly in South door sensor.

22:15 Online Split requests UPH036 ---> KTEV2/tps3d2.

22:28 Beam back. DAQ back. But NM3S is off.

15075 22:30 Back in business! Run 15075 - BEAM_799 has begun.

24:00 End of an exciting EVEning shift.

KTeV Logbook XXIII, pages 8-9.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shifts

--
===============================================================================
Earl C. Swallow                           email:  earls@elmhurst.edu
Theophil W. Mueller Professor              (or)   earls@fnal.gov
   and Chairperson
Department of Physics                     Phone:  (630)-617-3577
Elmhurst College                          FAX:    (630)-617-3735
190 Prospect Avenue
Elmhurst, IL   60126-3296                 http://www.elmhurst.edu/~earls
===============================================================================
"Books are lighthouses erected in the dark sea of time."
                                --  from Gargoyles  (The Scrolls of Merlin)